OpenStudy (anonymous):

540 is what % of 600, would I take 0.0540 and multiply by 600?

OpenStudy (anonymous):

if 600 is 100% 540 is x 540 * 100 / 600 90%

OpenStudy (anonymous):

gosh why this is confusing to me thank you so much!!!

OpenStudy (anonymous):

start by translating in to math. 540 IS what percent of 600 \[540= \frac{P}{100}\times 600\] \[\frac{540}{600}\times 100=P\] \[

OpenStudy (anonymous):

so 594 is 10% of what #? how ould I start this?

OpenStudy (anonymous):

same way. replace the "is" by an equal sign, and the "of" by a times sign, and the "what" by a variable or a letter, or just the word "what". try that and see if you understand.

OpenStudy (anonymous):

what and # go together right?

OpenStudy (anonymous):

yes. "what number" is what you are looking for. you can call it P or x or whatever.

OpenStudy (anonymous):

you always lead me to the light. thanx

OpenStudy (anonymous):

hey can one of you guys help me please :(

OpenStudy (anonymous):

ok did you write \[594=\frac{10}{100}\times x\]?

OpenStudy (anonymous):

in reality you just turn 10% to a decimal in your head. \[\frac{10}{100}=0.1\] so you write \[594=0.1\times x\] and solve for x

OpenStudy (anonymous):

59.4

OpenStudy (anonymous):

no my dear

OpenStudy (anonymous):

thats weird it was the original n umber but with a decimal in it

OpenStudy (anonymous):

you have to divide 594 by 0.1, not multiply 594 by 0.1

OpenStudy (anonymous):

nothing weird about that, any number times 1 is itself. but lets go slow ok?

OpenStudy (anonymous):

its really easy just use the rule of 3!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

in the tabbey.jpg add the fixed cost to the multiplication!

OpenStudy (anonymous):

i will listen because i don't know this rule. but if some number is ten percent of something, you have to multiply it by ten to get the something.

OpenStudy (anonymous):

xizzen hold up im almost done

OpenStudy (anonymous):

what is x?

OpenStudy (anonymous):

look! in this example. so 594 is 10% of what #? how ould I start this? with the rule of 3... it goes like this. if 10% is 594 100% is x 100 * 594 / 10 5940

OpenStudy (anonymous):

so it is 5940 ok I was right thank u I wanted to make clear